Show that the ideal <x^2+1> is not prime in Z2[x]
1
Expert's answer
2020-10-14T18:49:52-0400

An ideal "P" of a commutative ring "R" is prime if it has the following two properties:

1) If "a" and "b" are two elements of "R" such that their product "ab\in P", then "a\in P" or "b\in P";

2) "P" is not the whole ring "R".


Let "f(x)=x+1\in\mathbb Z_2[x]" and "g(x)=x+1\in\mathbb Z_2[x]". Since "1+1=0" in the ring "\mathbb Z_2", "f(x)g(x)=(x+1)(x+1)=x^2+x+x+1=x^2+(1+1)x+1=x^2+1" is in the ideal "I=\langle x^2+1\rangle" generated by "x^2+1." Taking into account that "f(x)\notin I" and "g(x)\notin I," we conclude that the ideal "I=\langle x^2+1\rangle" generated by "x^2+1" is not prime.